OpenStudy (anonymous):

Simplifying (fractions) Joaquim wants to make the macaroni and cheese shown to the right(2 1/2 cups of mc) but he only has about 1 3/4 cups of mc. About how much does he need?

OpenStudy (jhannybean):

\[\sf 1 \frac{3}{4} = \frac{4+3}{4} = \frac{7}{4}\]\[\sf 2 \frac{1}{2} = \frac{4+1}{2} = \frac{5}{2}\]Total mac n cheese = \(\sf \frac{5}{2}\) Has: \(\frac{7}{4}\) \[\sf \frac{7}{4} +x = \frac{5}{2}\]Solve for x. x is the amount of cheese he needs more of.

OpenStudy (anonymous):

kk thanks!

OpenStudy (jhannybean):

no problem :)